OpenStudy (anonymous):

The map shows the approximate location of some of Warwick important buildings. What building is located ate -1,0 I will give you the link of the map as soon as you chat with me :)

OpenStudy (anonymous):

I am listening to you :)

OpenStudy (anonymous):

I wil give you the link right now its right here :) thanks for helping :) the link name is super

OpenStudy (anonymous):

Police station. The coordinates are (-1,0) which means x = -1 and y = 0. At the point where x = -1 and y = 0 there is C, which is Police station.

OpenStudy (anonymous):

C is wear the police station is

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so the answer is police station?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you can you answer me a few more questions please?

OpenStudy (anonymous):

Which shows the correct line plot for the following data?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay above when you said yes is the link to another question

OpenStudy (anonymous):

I don't understand this question, what is tally and value? what are they refer to?

OpenStudy (anonymous):

sorry I took to long I was eating dinner any ways a tally is just lines to represent a number and a value is like how much its worth I think

OpenStudy (anonymous):

Hmm, ok I solved it but I don't know how to explain it to you.

OpenStudy (anonymous):

just give me the answer

OpenStudy (anonymous):

yea? just give me the answer

OpenStudy (anonymous):

The crosses are the frequencies (tallies and frequencies represent same thing btw). Before explaining I would like to show this: \[3\frac{1}{2} = 3 + \frac{1}{2}\] First lets look at the first one: \[3\frac{3}{4} = 3 + \frac{3}{4} = 3 + 0.75\] In the number line, the point 3.75 will have 2 crosses. Second: \[4\frac{1}{4} = 4 + \frac{1}{4} = 4 + 0.25 = 4.25\] In the number line, point 4.25 will get 5 crosses. You solve others like this. I hope I explained it well.

OpenStudy (anonymous):

The answer is the second choice